1a Differentiate the following functions.

Β 

𝑦=5π‘₯3βˆ’3π‘₯2+5π‘₯2βˆ’7 𝑀.π‘Ÿ.𝑑.π‘₯

Β 

𝑣=5𝑒2π‘¦βˆ’π‘’βˆ’3𝑦 𝑀.π‘Ÿ.𝑑.𝑦

Β 

𝑦=4𝑠𝑖𝑛2πœƒβˆ’2π‘π‘œπ‘ 4πœƒ 𝑀.π‘Ÿ.𝑑.πœƒ


1
Expert's answer
2021-04-14T07:27:10-0400



Solutions:-


(1) "\boxed{\frac{dy}{dx} = 15x^2 + 10x}"


(2) "\frac{dv}{dy} = 5e^2 (-e)y^{-e-1 \space }-3"


"\boxed{\frac{dv}{dy} = -5e^3y^{-e-1 \space }-3}"




(3) "\boxed{\frac{dy}{d \theta} = 8cos(2\theta) + 8sin(4\theta)}"